# A Survey of Text Watermarking in the Era of Large Language Models

### v2

@@ Abstract.
In recent years, significant advancements have been made in the text generation capabilities of Large Language Models (LLMs), demonstrating exceptional performance in downstream tasks such as abstract summarization, dialogue generation, and data-to-text conversion. However, their generative abilities also pose risks such as the rapid spread of fake news, infringement of datasets/LLM copyrights, and challenges to academic integrity. Text watermarking technology emerges as a potential solution. By embedding invisible yet detectable patterns in generated texts, it helps in tracking and verifying text origins, thus preventing misuse and piracy.
This survey aims to comprehensively summarize current text watermarking technologies, covering three main aspects: (1) an overview and comparison of different text watermarking techniques; (2) evaluation methods for text watermarking algorithms, including their success rate, impact on text quality, robustness, and unforgeability; (3) potential applications of text watermarking technologies. This survey aims to help researchers thoroughly understanding the text watermarking technologies, thereby fostering further development.
@@ 1. Introduction
However, the extensive use of LLMs has also introduced a set of problems and challenges. Firstly, the rapid generation of high-quality text by large models can facilitate the rapid spread of false information (Megías et al., 2021). Secondly, the issue of intellectual property related to large models is of vital importance. This includes the copyright of datasets (Tang et al., 2023) used for training large models and the addition of intellectual property rights to prevent the extraction of knowledge from the models (Zhao et al., 2023b). If effective tagging and detection methods for LLM-generated text could be implemented, it would significantly aid in mitigating the aforementioned issues.
@@
Text watermarking technology embeds specific identification information into the text data, which is difficult for humans to detect but could easily be identified by designed watermark detection algorithms, thereby serving as an effective solution to tag and detect LLM-generated text. Currently, various text watermarking methods have emerged, and these watermark algorithms are increasingly integrated with large language models. For instance, text generated by LLMs can be watermarked using text watermarking algorithms (Brassil et al., 1995; Por et al., 2012; Rizzo et al., 2016; Munyer and Zhong, 2023; Yang et al., 2022, 2023; Yoo et al., 2023b), or LLMs themselves can be utilized to embed watermarks in texts (Abdelnabi and Fritz, 2021; Zhang et al., 2023a). Additionally, watermark algorithms can be directly incorporated during the text generation process of LLMs (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023c, b; Ren et al., 2023; Wu et al., 2023). However, up to date, no studies have attempted to comprehensively explore and understand text watermarking from a broader perspective. The current surveys on text watermarking mainly focus on techniques prior to the era of large language models (Alkawaz et al., 2016; Kamaruddin et al., 2018).
@@
Therefore, in this work, we provide the first comprehensive survey of text watermarking algorithms in the era of large language models, covering the detailed definition of text watermarking algorithms and the interconnections between different kinds of text watermarking methods. Given the complexity and diversity of text watermarking technology, we have also detailed how to evaluate text watermarking algorithms from different perspectives, including success rate, robustness, impact on text quality and unforgeability. Additionally, we have introduced the current application scenarios of text watermarking algorithms including copyright protection, fake news detection, and academic integrity. This survey can provide researchers with a high-level understanding of text watermarking algorithms and a comparison of the similarities and differences between different text watermarking algorithms. Researchers who merely wish to employ text watermarking technology can also select the appropriate algorithms and application scenarios based on the introduction provided in this survey.
@@
Organization of this survey. The remainder of this survey is organized as follows: Section 2 introduces the definition of text watermarking and the essential properties of text watermarking algorithms. Section 3 and Section 4 discuss two significant types of text watermarking algorithms: text watermarking for existing text and text watermarking for LLMs. Section 5 elaborates on the different evaluation perspectives of text watermarking algorithms, including success rate, impact on text quality, robustness, and unforgeability. Section 6 presents the application scenarios of text watermarking algorithms, covering copyright protection, academic integrity and fake news detection. Finally, Section 7 concludes the survey.
@@ 2.3. Taxonomy of Text Watermarking Algorithms
Text Watermarking Watermarking for Existing Text (section˜3) Format-based Watermarking (section˜3.1) Line/Word-Shift Coding (Brassil et al. (1995)), UniSpaCh (Por et al. (2012)), Unicode Homoglyph Substitution (Rizzo et al. (2016)), Embarrassingly Simple Watermark (Sato et al. (2023)) Lexical-based Watermarking (section˜3.2) Synonym Substitutions (Topkara et al. (2006b)), Deeptextmark (Munyer and Zhong (2023)), Context-aware Lexical Substitution (Yang et al. (2022)), Binary-encoding Lexical Substitution (Yang et al. (2023)), Robust Multi-bit Watermark (Yoo et al. (2023b)) Syntatic-based Watermarking (section˜3.3) NLW (Atallah et al. (2001)), WANE (Topkara et al. (2006a)), MA-NLW (Meral et al. (2009)) Generation-based Watermarking ( section˜3.4) AWT (Abdelnabi and Fritz (2021)), REMARK-LLM (Zhang et al. (2023a)) Watermarking for LLMs (section˜4) Training Time Watermarking (section˜4.1) Dataset Trigger (Liu et al. (2023a)), Clean-Label Backdoor Watermark (Tang et al. (2023)), Coprotector (Sun et al. (2022)), CodeMark (Sun et al. (2023)) Watermarking During Logits Generation (section˜4.2) KGW (Kirchenbauer et al. (2023b)), SWEET (Lee et al. (2023)), Unbiased Watermark (Hu et al. (2023)), DiPmark (Wu et al. (2023)), MWPA (Yoo et al. (2023a)), CTWL (Wang et al. (2023b)), Unigram Watermark (Zhao et al. (2023a)), KGW-reliability (Kirchenbauer et al. (2023c)), Semantic Invariant Robust Watermark (Liu et al. (2023c)), Unforgeable Watermark (Liu et al. (2023b)), Public Detectable Watermark (Fairoze et al. (2023a)), Semantic-based Robust Watermark (Ren et al. (2023)) Watermarking During Token Sampling (section˜4.3) Undetectable Watermark (Christ et al. (2023)), Robust Distortion-free Watermark (Kuditipudi et al. (2023))
@@ 3.1. Format-based Watermarking
Format-based watermarking approaches are inspired by image watermarking technology (Begum and Uddin, 2020). It does not modify the content of the text but introduces changes to its format that are difficult for humans to detect, thereby embedding a watermark. For example, Brassil et al. (1995) proposed line-shift coding and word shift-coding techniques, achieved by vertically shifting the positions of text lines or horizontally shifting the locations of words within text lines. Correspondingly, the watermark detection process involves measuring the distance between adjacent text line profiles or between adjacent word column profiles to detect shifts. However, this approach is limited to embedding watermarks in image-formatted text and cannot truly return a text string with an embedded watermark. Considering this, various watermarking methods that rely on the insertion or replacement of Unicode codepoints have been introduced. Por et al. (2012) proposed a watermarking scheme named UniSpach, which inserts Unicode space characters into inter-sentence, inter-word, end-of-line and inter-paragraph spacings. A study by Rizzo et al. (2016) presented a unicode homoglyph substitution text watermarking method. It exploits the fact that text symbols that are similar in appearance could have different Unicode codepoints. For instance, both U+0043 and U+216d visually represent letter ‘C’, while U+004c and U+216c appear as letter ‘L’. Following this, a family of simple watermarks named EASYMARK (Sato et al., 2023) is proposed recently, composed of three different methods: WHITEMARK, VARIANTMARK and PRINTMARK. Specifically, WHITEMARK replaces a whitespace (U+0020) with another codepoint of a whitespace (e.g. U+2004). VARIANTMARK leverages variation selectors of Unicode to embed hard-to-perceive format into CJK texts. PRINTMARK, coping with printed texts, uses ligature or whitespaces with slightly different lengths to embed watermark messages. As these watermarking methods relied on the richness of Unicode encoding, the watermark detection process involves searching for and counting the certain codepoints that have been inserted within the text.
However, although these format-based watermarking methods allow for the simple and effective embedding of watermarks in text without altering its specific content, modifications to the format can be easily spotted in certain scenarios, as mentioned by Por et al. (2012) in the case of DASH attack. As a result, these specially designed formats could be effortlessly removed through canonicalization (Boucher et al., 2022), such as reseting line spacing, searching and replacing certain codepoints through the entire text, etc. The spotted formats might also be used to forge watermark, further leading to the failure of effective detection.
@@ 3.2. Lexical-based Watermarking
Format-based approaches only modify the surface format of the text, making them easily spotted and, consequently, more vulnerable to targeted removal by reformatting. Therefore, it becomes imperative to explore alternative methods that enable deeper insertion of watermarks within text. Several studies applied word-level modifications by replacing selected words with their alternatives without changing the sentence syntax structure. Topkara et al. (2006b); Fellbaum (1998); Munyer and Zhong (2023); Yang et al. (2022, 2023); Yoo et al. (2023b) We refer to these methods as lexical-based watermarking approaches. Topkara et al. (2006b) presented a synonym substitution text watermarking approach, using the linguistic database WordNet(Fellbaum, 1998) as its synonym dictionary. The watermark detection process essentially replicates the watermark message embedding procedure, with the distinction that inverse rules are employed during the message extraction phase. In order to conduct semantic modeling more effectively, Munyer and Zhong (2023) utilized a pretrained Word2Vec model instead of WordNet to find alternatives. In particular, they converted the selected words into Word2Vec vectors and collected the n-nearest vectors to form replacement word set. Noticeably, they trained a binary classifier as watermark detector, using pretrained BERT model and transformer blocks as neural network components.
@@ 4. Watermarking for LLMs
To facilitate explanation, we assume that the watermark text is directly generated by a language language model $M_{w}$ with an embedded watermark message.
@@ 4.2. Watermarking during Logits Generation
Kirchenbauer et al. (2023a) proposed the first watermarking method based on LLM logits modification, referred to as KGW. There, the watermark generator randomly splits the vocabulary set into a red list and a green list at each position depending on the previous token using a hash function. As illustrated in the left part of Figure 3, when $M_{w}$ generates the $i^{th}$ token, a small bias $\delta$ is added to the logits of the tokens in the green list. Let $G$ represent the green list and $R$ stands for the red list. Then the logits value of token $v_{j}$ at position $i$ would be calculated as follows:
@@
Figure 3. A more illustrative description of watermarking during logits generation: The left fugure presents the classic KGW (Kirchenbauer et al., 2023a) algorithm in this field, while the right fugure demonstrates how various methods improve upon this approach to adapt to four different scenarios: Watermarking Low-entropy Text, Watermark with Multi-bit Payload, Preventing Watermark Removal Attack, and Defending against Watermark Forgeries.
Although the KGW watermark detection method demonstrated exceptional performance in its test scenarios, achieving a false positive rate (human text misclassified as watermark text) of less than $3\times 10^{-3}$%, and a false negative rate (watermark text not detected) of less than 1%, there remain numerous highly challenging scenarios in real-world applications that necessitate specialized optimization and design of the watermark algorithm to effectively address them. Below, we list four representative scenarios and thoroughly introduce the improvements and explorations made to watermark algorithms in these contexts, which are illustrated in the right part of Figure 3.
@@
The watermark algorithm introduced by Equation 9, known as the KGW(Kirchenbauer et al., 2023a), could only determine the presence of a watermark and no additional information, categorizing it as a zero-bit payload watermark. However, practical applications often demand that watermarks carry additional information such as copyright details, timestamps, or specific identifiers. This necessitates watermarking techniques that not only detect the presence of watermarks but also extract meaningful information, which could be categorized as multi-bit payload watermark.
To achieve this, one possible solution is to establish multiple splits for dividing the vocabulary into a red list and a green list. Specifically, we can set N types of splits for the vocabulary, such as $(G_{1},R_{1}),…,(G_{N},R_{N})]$, where each split could performing LLM watermarking following the way introduced in Equation 9. Each split corresponds to a specific watermark message, endowing the watermark with a $log_{2}{N}$-bit payload. In this context, various approaches are adopted to design the split method for a vocabulary differently. For instance, Wang et al. (2023b) permits the user to input a watermark message, $m$. Subsequently, based on the hash value of this message $m$, the vocabulary is divided. Given that the hash values vary with different messages, this methodology effectively achieves distinct vocabulary splitting techniques for each unique message. However, this method requires multiple times of detection to extract the multi-bit payload and is thus not efficient.
To enable more efficient detection of the multi-bit payload watermark, Yoo et al. (2023a) expanded the splitting rules from red-green partition to multi-color partition. Different watermark messages result in preference towards the tokens of different colors during logits generation process. During watermark detection phrase, this approach only requires counting the number of tokens in the subsets of vocabulary corresponding to different colors in the text thus achieving more efficient detection. However, in the case of multi-bit payload watermark algorithms, while they enhance the watermark payload, they also tend to result in a reduction in the quality of the generated text or lower the robustness of watermark, meaning that modifications to the text more easily lead to changes in the watermark message. Future work still needs to balance the trade-offs among the watermark’s payload, robustness, and its impact on text quality.
@@
In particular, for the watermarking during logits generation methods (Kirchenbauer et al., 2023a, c; Zhao et al., 2023a; Liu et al., 2023b, c), the most significant factor affecting robustness is how to determine the modifications to logits. More specifically, how to divide the red-green list mentioned in the KGW method (Kirchenbauer et al., 2023a). In the original KGW method (Kirchenbauer et al., 2023a), the red-green list is determined based on the hash value of preceding tokens. Kirchenbauer et al. (2023c) further elaborated on some specific hash strategies, such as using only the token with the smallest token id in previous tokens for hashing to decide the red-green list, which can enhance robustness. (Zhao et al., 2023a) proved that a globally fixed split of red and green list results in higher robustness to watermark removal attacks . Additionally, since watermark removal attacks typically do not alter the semantic content of the text, many studies have also designed methods to determine the split of the red-green list based on textual semantics. For example, Liu et al. (2023c) trained a watermark model that can directly convert text semantic embeddings into watermark logits. Ren et al. (2023) converted semantic embeddings into semantic values through weighted embedding pooling followed by discretizing using NE-Ring, and then divided the vocabulary into red-list and green-list based on these semantic values. The current methods primarily focus on investigating the robustness of zero-bit watermark algorithms against watermark removal attacks. Future algorithms could further explore the robustness of multi-bit payload watermark algorithms.
@@
In the aforementioned discussion on the watermark algorithms’ ability to preventing watermark removal attacks, it is assumed that the attacker is not aware of the details and generation methods of the watermark algorithm, which is essentially a black-box setting. Once an attacker obtain the watermark generation details, they could effortlessly remove the watermark using the anti-watermark techniques (Kirchenbauer et al., 2023c) or forge the watermark easily. Therefore, for watermark algorithms, the capability to defend against watermark forgeries is of great importance. The ability to defend against watermark forgeries depends on the watermark algorithm’s capacity to effectively conceal its watermark generation process.
@@
In the context of public detection scenarios, resisting watermark forgeries becomes significantly more challenging. This difficulty arises because attackers can access the watermark detectors. For methods where the watermark generation process is involved in detection (Kirchenbauer et al., 2023a, c; Zhao et al., 2023a; Liu et al., 2023c), exposing the watermark generator leads to a complete inability to resist watermark forgeries. To address this issue, Fairoze et al. (2023b) have utilized digital signature technology from the field of cryptography. This approach involves generating watermarks using a private key and verifying them with a public key. However, the verification via public key relies on features extracted from the text and users can still exploit these features to forge watermarks. Further advancing this field, Liu et al. (2023b) proposed the use of neural networks for watermark detection. Due to the black-box nature of neural networks, the details of watermark generation are not exposed, which could defend against watermark forgeries in public detection scenarios.

### v3

@@ Abstract.
Text watermarking algorithms play a crucial role in the copyright protection of textual content, yet their capabilities and application scenarios have been limited historically. The recent developments in large language models (LLMs) have opened new opportunities for the advancement of text watermarking techniques. LLMs not only enhance the capabilities of text watermarking algorithms through their text understanding and generation abilities but also necessitate the use of text watermarking algorithms for their own copyright protection. This paper conducts a comprehensive survey of the current state of text watermarking technology, covering four main aspects: (1) an overview and comparison of different text watermarking techniques; (2) evaluation methods for text watermarking algorithms, including their success rates, impact on text quality, robustness, and unforgeability; (3) potential application scenarios for text watermarking technology; (4) current challenges and future directions for development. This survey aims to provide researchers with a thorough understanding of text watermarking technology, thereby promoting its further advancement.
@@ 1. Introduction
Text watermarking involves embedding unique, imperceptible identifiers (watermarks) into textual content. These watermarks are designed to be robust yet inconspicuous, ensuring that the integrity and ownership of the content are preserved without affecting its readability or meaning. Historically, text watermarking has played a crucial role in various domains, from copyright protection and document authentication to preventing plagiarism and unauthorized content distribution (Kamaruddin et al., 2018). With the advancement of Large Language Models (LLMs), both the techniques and application scenarios of text watermarking have seen significant development. As shown in Figure 1(a), this primarily includes the construction of enhanced text watermarking algorithms using LLMs, the application of existing text watermarking algorithms to LLMs, and the exploration of text watermarking algorithms that are more closely integrated with LLMs. The flourishing development of LLMs has propelled a thriving research landscape within the realm of text watermarking, as depicted in Figure 1(b). Especially with the advent of ChatGPT, text watermarking has notably surged into a research fervor. To help better understanding the mutually beneficial relationship between LLMs and text watermarking, this paper provides a survey of text watermarking techniques in the era of large language models.
In the subsequent content of this section, we will separately discuss why text watermarking benefits the application of LLMs (section 1.1), why utilizing LLMs can lead to the development of superior text watermarking algorithms (section 1.2), and the contributions of this survey along with the organization of the following sections (section 1.3).
(a) A description of how LLMs promote the development of text watermarking techniques and broaden their application scenarios.
(b) Number of publications in the field of text watermarking and LLMs (the data for "Number of Publications in the field of LLMs" is sourced from Zhao et al. (2023c))
Figure 1. Relationships between the development of text watermarking techniques and Large Language Models (LLMs).
1.1. Why is Text Watermarking Beneficial for LLMs?
@@
However, the extensive use of LLMs has also introduced a set of problems and challenges. Firstly, the rapid generation of high-quality text by LLMs can facilitate the rapid spread of false information (Megías et al., 2021). Secondly, the issue of intellectual property related to large models is of vital importance. This includes the copyright of datasets (Tang et al., 2023) used for training large models and the addition of intellectual property rights to prevent the extraction of knowledge from the models (Zhao et al., 2023b). If effective tagging and detection methods for LLM-generated text could be implemented, it would significantly aid in mitigating the aforementioned issues. Text watermarking emerges as a promising solution to address these challenges. By embedding a unique, identifiable, and non-obtrusive marker (watermark) within the LLM-generated text, watermarking can enable the tracking and attribution of content produced by LLMs.
1.2. Why are LLMs Beneficial for Text Watermarking?
One of the main challenges in text watermarking is embedding watermarks without altering the original meaning or readability of the text. This requirement presents a significant challenge, as traditional text watermarking methods often struggle to modify the text without changing its semantics (Atallah et al., 2001; Topkara et al., 2006a; Meral et al., 2009). This is due to the need for text watermarking algorithms to have a strong understanding and control over the text’s semantics. Here, Large Language Models emerge as a game-changer. With their advanced understanding of language semantics and context, LLMs enable more sophisticated text watermarking methods that seamlessly integrate watermarks with minimal compromise to the text’s original meaning (Abdelnabi and Fritz, 2021; Zhang et al., 2023a). This synergy allows for the development of watermarking techniques that are not only more effective but also more subtle, ensuring that the text remains as intended while still carrying the necessary watermark features.
1.3. Why a Survey for Text Watermarking in the Era of LLMs?
@@
Although text watermarking technology and large language models can effectively enhance each other, for instance, text generated by LLMs can be watermarked using text watermarking algorithms (Brassil et al., 1995; Por et al., 2012; Rizzo et al., 2016; Munyer and Zhong, 2023; Yang et al., 2022, 2023; Yoo et al., 2023b), or LLMs themselves can be utilized to embed watermarks in texts (Abdelnabi and Fritz, 2021; Zhang et al., 2023a). Additionally, watermark algorithms can be directly incorporated during the text generation process of LLMs (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023c, b; Ren et al., 2023; Wu et al., 2023). However, up to date, no studies have attempted to comprehensively explore and understand text watermarking from a broader perspective. The current surveys on text watermarking mainly focus on techniques prior to the era of large language models (Alkawaz et al., 2016; Kamaruddin et al., 2018).
@@
Therefore, in this work, we provide the first comprehensive survey of text watermarking algorithms in the era of large language models, covering the detailed definition of text watermarking algorithms and the interconnections between different kinds of text watermarking methods. Given the complexity and diversity of text watermarking technology, we have also detailed how to evaluate text watermarking algorithms from different perspectives, including success rate, robustness, impact on text quality and unforgeability. Additionally, we have introduced the current application scenarios of text watermarking algorithms including copyright protection, fake news detection, and academic integrity. This survey can provide researchers with a high-level understanding of text watermarking algorithms and a comparison of the similarities and differences between different text watermarking algorithms. Researchers who merely wish to employ text watermarking technology can select the appropriate algorithms and application scenarios based on the introduction provided in this survey.
@@
Organization of this survey. The remainder of this survey is organized as follows: Section 2 introduces the definition of text watermarking and the essential properties of text watermarking algorithms. Section 3 and Section 4 discuss two significant types of text watermarking algorithms: text watermarking for existing text and text watermarking for LLMs. Section 5 elaborates on the different evaluation perspectives of text watermarking algorithms, including success rate, impact on text quality, robustness, and unforgeability. Section 6 presents the application scenarios of text watermarking algorithms, covering copyright protection, academic integrity and fake news detection. Section 7 analyzes the challenges still faced by current text watermarking research and explores future research directions. Finally, Section 8 concludes the survey.
@@ 2.3. Taxonomy of Text Watermarking Algorithms
Text Watermarking Watermarking for Existing Text (section˜3) Format-based Watermarking (section˜3.1) Line/Word-Shift Coding (Brassil et al. (1995)), UniSpaCh (Por et al. (2012)), Unicode Homoglyph Substitution (Rizzo et al. (2016)), EasyMark (Sato et al. (2023)) Lexical-based Watermarking (section˜3.2) Equimark (Topkara et al. (2006b)), DeepTextMark (Munyer and Zhong (2023)), Context-aware Lexical Substitution (Yang et al. (2022)), Binary-encoding Lexical Substitution (Yang et al. (2023)), Robust Multi-bit Watermark (Yoo et al. (2023b)) Syntatic-based Watermarking (section˜3.3) NLW (Atallah et al. (2001)), WANE (Topkara et al. (2006a)), MA-NLW (Meral et al. (2009)) Generation-based Watermarking ( section˜3.4) AWT (Abdelnabi and Fritz (2021)), REMARK-LLM (Zhang et al. (2023a)) Watermarking for LLMs (section˜4) Training Time Watermarking (section˜4.1) Dataset Trigger (Liu et al. (2023a)), Clean-Label Backdoor Watermark (Tang et al. (2023)), Coprotector (Sun et al. (2022)), CodeMark (Sun et al. (2023)) Watermarking During Logits Generation (section˜4.2) KGW (Kirchenbauer et al. (2023a)), SWEET (Lee et al. (2023)), Unbiased Watermark (Hu et al. (2023)), DiPmark (Wu et al. (2023)), COLOR (Yoo et al. (2023a)), CTWL (Wang et al. (2023b)), ThreeBricks (Fernandez et al. (2023)), Unigram Watermark (Zhao et al. (2023a)), KGW-reliability (Kirchenbauer et al. (2023b)), NS-Watermark (Takezawa et al. (2023)), Semantic Invariant Robust Watermark (Liu et al. (2023c)), Unforgeable Watermark (Liu et al. (2023b)), Publicly Detectable Watermark (Fairoze et al. (2023)), Semantic-based Robust Watermark (Ren et al. (2023)) Watermarking During Token Sampling (section˜4.3) Undetectable Watermark (Christ et al. (2023)), Robust Distortion-free Watermark (Kuditipudi et al. (2023)), SemStamp (Hou et al. (2023)))
@@ 3.1. Format-based Watermarking
Format-based watermarking approaches are inspired by image watermarking technology (Begum and Uddin, 2020). It does not modify the content of the text but introduces changes to its format that are difficult for humans to detect, thereby embedding a watermark. For example, Brassil et al. (1995) proposed line-shift coding and word shift-coding techniques, achieved by vertically shifting the positions of text lines or horizontally shifting the locations of words within text lines. Correspondingly, the watermark detection process involves measuring the distance between adjacent text line profiles or between adjacent word column profiles to detect shifts. However, this approach is limited to embedding watermarks in image-formatted text and cannot truly return a text string with an embedded watermark. Considering this, various watermarking methods that rely on the insertion or replacement of Unicode codepoints have been introduced. Por et al. (2012) proposed a watermarking scheme named UniSpach, which inserts Unicode space characters into inter-sentence, inter-word, end-of-line and inter-paragraph spacings. A study by Rizzo et al. (2016) presented a unicode homoglyph substitution text watermarking method. It exploits the fact that text symbols that are similar in appearance could have different Unicode codepoints. For instance, both U+0043 and U+216d visually represent letter ‘C’, while U+004c and U+216c appear as letter ‘L’. Following this, a family of simple watermarks named EASYMARK (Sato et al., 2023) is proposed recently, composed of three different methods: WHITEMARK, VARIANTMARK and PRINTMARK. Specifically, WHITEMARK replaces a whitespace (U+0020) with another codepoint of a whitespace (e.g. U+2004). VARIANTMARK leverages variation selectors of Unicode to embed hard-to-perceive format into CJK texts. PRINTMARK, coping with printed texts, uses ligature or whitespaces with slightly different lengths to embed watermark messages. Correspondingly, the watermark detection process involves searching for and counting the certain codepoints that have been inserted within the text. As these watermarking methods relied on the richness of Unicode encoding, their watermark payload are often quite large.
However, although these format-based watermarking methods allow for the simple and effective embedding of large payload watermarks in text without altering its specific content, modifications to the format can be easily spotted in certain scenarios, as mentioned by Por et al. (2012) in the case of DASH attack. As a result, these specially designed formats could be effortlessly removed through canonicalization (Boucher et al., 2022), such as reseting line spacing, searching and replacing certain codepoints through the entire text, etc. The spotted formats might also be used to forge watermark, further leading to the failure of effective detection.
@@ 3.2. Lexical-based Watermarking
Format-based approaches only modify the surface format of the text, making them easily spotted and, consequently, more vulnerable to targeted removal by reformatting. Therefore, it becomes imperative to explore alternative methods that enable deeper insertion of watermarks within text. Several studies applied word-level modifications by replacing selected words with their alternatives without changing the sentence syntax structure (Topkara et al., 2006b; Fellbaum, 1998; Munyer and Zhong, 2023; Yang et al., 2022, 2023; Yoo et al., 2023b). We refer to these methods as lexical-based watermarking approaches. Topkara et al. (2006b) presented a synonym substitution text watermarking approach, using the linguistic database WordNet(Fellbaum, 1998) as its synonym dictionary. The watermark detection process essentially replicates the watermark message embedding procedure, with the distinction that inverse rules are employed during the message extraction phase. In order to conduct semantic modeling more effectively, Munyer and Zhong (2023) utilized a pretrained Word2Vec model instead of WordNet to find alternatives. In particular, they converted the selected words into Word2Vec vectors and collected the n-nearest vectors to form replacement word set. Noticeably, they trained a binary classifier as watermark detector, using pretrained BERT model and transformer blocks as neural network components.
@@ 4. Watermarking for LLMs
To facilitate explanation, we assume that the watermarked text is directly generated by a language language model $M_{w}$ with an embedded watermark message.
@@ 4.2. Watermarking during Logits Generation
Kirchenbauer et al. (2023a) proposed the first watermarking method based on LLM logits modification, referred to as KGW. There, the watermark generator randomly splits the vocabulary set into a red list and a green list at each position depending on the previous token using a hash function. As illustrated in Figure 4, when $M_{w}$ generates the $i^{th}$ token, a small bias $\delta$ is added to the logits of the tokens in the green list. Let $G$ represent the green list and $R$ stands for the red list. Then the logits value of token $v_{j}$ at position $i$ would be calculated as follows:
@@
Figure 4. A more illustrative description of the KGW (Kirchenbauer et al., 2023a) algorithm.
Although the KGW watermark detection method demonstrated exceptional performance in its test scenarios, achieving a false positive rate (human text misclassified as watermark text) of less than $3\times 10^{-3}$%, and a false negative rate (watermark text not detected) of less than 1%, there remain numerous highly challenging scenarios in real-world applications that necessitate specialized optimization and design of the watermark algorithm to effectively address them. Below, we list four representative scenarios and thoroughly introduce the improvements and explorations made to watermark algorithms in these contexts, which are illustrated in Figure 5.
@@
The KGW(Kirchenbauer et al., 2023a) watermark algorithm introduced by Equation 9, could only determine the presence of a watermark and no additional information, categorizing it as a zero-bit payload watermark. However, practical applications often demand that watermarks carry additional information such as copyright details, timestamps, or specific identifiers. This necessitates watermarking techniques that not only detect the presence of watermarks but also extract meaningful information, which could be categorized as multi-bit payload watermark.
To achieve this, one possible solution is to establish multiple splits for dividing the vocabulary into a red list and a green list. Specifically, we can set N types of splits for the vocabulary, such as $(G_{1},R_{1}),…,(G_{N},R_{N})]$, where each split could performing LLM watermarking following the way introduced in Equation 9. Each split corresponds to a specific watermark message, endowing the watermark with a $log_{2}{N}$-bit payload. For instance, Wang et al. (2023b) permits the user to input a watermark message $m$ with $log_{2}{N}$ bits. Subsequently, based on the hash value of this message $m$, the vocabulary is divided. However, during detection phase, this method requires iterating through all $N$ possible messages, calculating the correlation between the text and the red/green splits corresponding to each message respectively, thus is not efficient. To mitigate this issue, Fernandez et al. (2023) proposed the Cyclic Shift method, where distinct messages are generated by cyclically shifting an initial message. This approach reduces redundant computational steps and enhances detection efficiency through parallel processing.
However, these methods, due to the necessity of iterating each possible message, encounter heightened computational complexity with the growing watermark payload. To address this, Yoo et al. (2023a) proposed allocating different bits of the message to various positions in the watermark text, allowing independent detection of different bit altogether in a single detection round. Then the information corresponding to each bit is concatenated to obtain the final message. Specifically, the message is modeled as a sequence $m=\Sigma^{b}$ where $\Sigma=\{0,1\}$. The process of dividing the vocabulary at each generation step is then split into two phrases: the first involves choosing a position in the message, $\Sigma^{b}[i]$, based on the existing random seed, and the second involves dividing the vocabulary according to this $\Sigma^{b}[i]$ and random seed. This approach significantly improves efficiency by enabling the parallel detection of bit information. Furthermore, Yoo et al. (2023a) also suggest dividing the vocabulary into more parts, so that each position $\Sigma^{b}[i]$ in $m$ can contain more information, further enhancing the payload of the text watermark.
@@
In particular, for the watermarking during logits generation methods (Kirchenbauer et al., 2023a, b; Zhao et al., 2023a; Liu et al., 2023b, c), the most significant factor affecting robustness is how to determine the modifications to logits. More specifically, how to divide the red-green list mentioned in the KGW method (Kirchenbauer et al., 2023a). In the original KGW method (Kirchenbauer et al., 2023a), the red-green list is determined based on the hash value of preceding tokens. Kirchenbauer et al. (2023b) further elaborated on some specific hash strategies, such as using only the token with the smallest token id in previous tokens for hashing to decide the red-green list, which can enhance robustness. (Zhao et al., 2023a) proved that a globally fixed split of red and green list results in higher robustness to watermark removal attacks . Additionally, since watermark removal attacks typically do not alter the semantic content of the text, many studies have also designed methods to determine the split of the red-green list based on textual semantics. For example, Liu et al. (2023c) trained a watermark model that can directly convert text semantic embeddings into watermark logits. Ren et al. (2023) converted semantic embeddings into semantic values through weighted embedding pooling followed by discretizing using NE-Ring, and then divided the vocabulary into red-list and green-list based on these semantic values. The current methods primarily focus on investigating the robustness of zero-bit watermark algorithms against watermark removal attacks. Future algorithms could further explore the robustness of multi-bit payload watermark algorithms.
Figure 5. Demonstration of how various methods improve upon KGW (Kirchenbauer et al., 2023a) to adapt to four scenarios: Watermarking Low-entropy Text, Watermark with Multi-bit Payload, Preventing Watermark Removal Attack, and Defending against Watermark Forgeries.
@@
In the aforementioned discussion on the watermark algorithms’ ability to preventing watermark removal attacks, it is assumed that the attacker is not aware of the details and generation methods of the watermark algorithm, which is essentially a black-box setting. Once an attacker obtain the watermark generation details, they could effortlessly remove the watermark using the anti-watermark techniques (Kirchenbauer et al., 2023b) or forge the watermark easily. Therefore, for watermark algorithms, the capability to defend against watermark forgeries is of great importance. The ability to defend against watermark forgeries depends on the watermark algorithm’s capacity to effectively conceal its watermark generation process.
@@
In the context of public detection scenarios, resisting watermark forgeries becomes significantly more challenging. This difficulty arises because attackers can access the watermark detectors. For methods where the watermark generation process is involved in detection (Kirchenbauer et al., 2023a, b; Zhao et al., 2023a; Liu et al., 2023c), exposing the watermark generator leads to a complete inability to resist watermark forgeries. To address this issue, Fairoze et al. (2023) have utilized digital signature technology from the field of cryptography. This approach involves generating watermarks using a private key and verifying them with a public key. However, the verification via public key relies on features extracted from the text and users can still exploit these features to forge watermarks. Further advancing this field, Liu et al. (2023b) proposed the use of neural networks for watermark detection. Due to the black-box nature of neural networks, the details of watermark generation are not exposed, which could defend against watermark forgeries in public detection scenarios.